Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-21T16:29:48-0400

.ANSWER.

1)An open interval is a set "(a,b)=\left\{ x\in R:a<x<b \right\}" .Let "c\in (a,b)" and "\varepsilon =min\left\{ \frac { b-c }{ 2 } ,\frac { c-a }{ 2 } \right\}" ."\varepsilon <\frac { b-c }{ 2 }" and "-ε>\frac { -c+a }{ 2 }" . Therefore

"a=\frac { a+a }{ 2 } <\frac { c+a }{ 2 } =\frac { 2c+a-c }{ 2 } =c+\frac { a-c }{ 2 } \ <c-\varepsilon \ <c+\varepsilon <c+\frac { b-c }{ 2 } =\frac { b+c }{ 2 } <\frac { 2b }{ 2 } =b" .

So, for all "c\in (a,b)\quad" exist ε>0 , such that "(c-\varepsilon \ ,c+\varepsilon )\subset (a,b)" .By the definition of an open set , any open interval is an open set.

2)Similarly, the sets "(-\infty ,a)=\left\{ x\in R:-\infty <x<a \right\}" and "(b,+\infty )==\left\{ x\in R:b<x<+\infty \right\}" are open sets . Because , for all "c\in \ (-\infty ,a)" exist "\varepsilon =\frac { -c+a }{ 2 }","(c+\varepsilon =\frac { c+a }{ 2 } <\frac { a+a }{ 2 } =a)" , such that "(c-\varepsilon \ ,c+\varepsilon )\subset (-\infty ,a)" .

If "d\in (b,+\infty )\quad \varepsilon =\frac { d-b }{ 2 } \quad (d-\varepsilon =\frac { d+b\quad }{ 2 } >b)" , then "(d-\varepsilon \quad ,d+\varepsilon )\subset (b,+\infty \quad )" .

3) A closed interval is a set "[a,b]=\left\{ x\in R:a\le x\ \le b \right\}" "=R\diagdown \left( (-\infty ,a)\cup (b,+\infty ) \right)"

That is, the closed interval is the complement of the set "I=(-\infty ,a)\cup (b,+\infty )" .From 2) it follows that the set "I" is open. Therefore, by the definition of the closed set, the set

"[a,b]" is closed.